OpenStudy (anonymous):

2x^2+8x+3=0?? complete the square on. leave in the from a(x+b)^2+c.

OpenStudy (anonymous):

I am stuck on 2[(x+2)^2-4+2/3]=0

OpenStudy (jamesj):

Because you have a 2 on the outside of the square brackets, and we're trying to get this in the form: a(x+b)^2+c

OpenStudy (anonymous):

I don't understand, sorry.

OpenStudy (anonymous):

But If I have to distribute the 2, should we also multiply the (x+2)^2 by 2 as well?

OpenStudy (anonymous):

you did you multiply 2(x+2)^2 by 2??

OpenStudy (anonymous):

it just stayed the same throughout.

OpenStudy (jamesj):

x^2 + 4x + 3/2 = x^2 + 4x + 4 - 4 + 3/2 = (x +2)^2 -4 + 3/2, because (x+2)^2 = x^2 + 4x + 4 Now does it make sense?

OpenStudy (jamesj):

Given what I've written above, it follows that 2[ x^2 + 4x + 3/2 ] = 2[ (x+2)^2 + (-4+3/2) ] that is, line A gives line B

OpenStudy (anonymous):

I am going to play around with it. thank you for your help.

OpenStudy (jamesj):

ok